# Robust and Real-Time Bangladeshi Currency Recognition: A Dual-Stream MobileNet and EfficientNet Approach

### v1

@@ I Introduction
  • We have proposed a novel hybrid model that combines MobileNetV3-Large and EfficientNetB0 as complementary feature extractors, followed by a multilayer perceptron (MLP) pipeline for classification.
@@ III-A1 Data Collection
Data collection is the first and fundamental step of research. To build our Bangladeshi banknote datasets, we have collected images of the currently available and frequently used Bangladeshi currency in 9 denominations (2 Tk, 5 Tk, 10 Tk, 20 Tk, 50 Tk, 100 Tk, 200 Tk, 500 Tk, and 1000 Tk). The images were collected by taking pictures from a mobile camera, a laptop webcam, and from publicly available datasets [29], [30], [40]. Figure 2 shows representative samples from the datasets that illustrate the progressive increase in visual variability and environmental complexity.
(a) Custom Dataset-1
(b) Primary Dataset-1
(c) Custom Dataset-2
(d) Primary Dataset-2
(e) Custom Dataset-3
Formally, the complete dataset $D$ can be expressed as the union of the sources:
@@ III-A2 Dataset Description
III-B Data Preprocessing and Augmentation
In order to preserve the discriminative feature integrity and improve model generalization, a structured preprocessing and augmentation pipeline was adopted. All the images were converted to JPEG format and resized to $224\times 224$ pixels, maintaining consistency and compatibility with the model.
III-B1 Data Preprocessing
The data preprocessing approach includes:
Normalization: Each image underwent pixel intensity normalization. The normalization process can be defined as:
$I_{\text{norm}}(x,y)=\frac{I(x,y)-\mu}{\sigma},$ (3)
where $I(x,y)$ denotes the pixel intensity at location $(x,y)$, and $\mu$ and $\sigma$ represent the mean and standard deviation, respectively. For transfer learning, we used the standard ImageNet normalization statistics. This centers and scales the data, which helps the model converge steadily during training.
Noise Reduction: To denoise the image but preserve the edges, the Non-Local Means (NLM) denoising algorithm (fastNlMeansDenoisingColored) was used with:
$I_{\text{denoise}}(p)=\frac{1}{C(p)}\sum_{q\in\Omega}w(p,q)\,I(q),$ (4)
where $\Omega$ is the search window, $w(p,q)$ is the patch similarity, and $C(p)$ is the normalization factor. This step cuts down on artifacts from lighting changes or camera inconsistencies.
Contrast Enhancement: CLAHE was used to boost the faded text or soft security details on the L-channel of the LAB color space:
$I_{\text{clahe}}(x,y)=\text{round}\left((L_{\text{max}}-L_{\text{min}})\cdot CDF(I(x,y))+L_{\text{min}}\right),$ (5)
where $CDF$ represents the cumulative distribution function of pixel intensities of a local area. The images were later converted back to RGB format to be compatible with the model.
III-B2 Data Augmentation
To enhance the model’s generalization ability in the real-world scenario, we did a series of augmentations on the training set on-the-fly via the Keras ImageDataGenerator where:
Rotation: Random rotations where $\theta\in[-30^{\circ},+30^{\circ}]$:
$\begin{bmatrix}x^{\prime}\\ y^{\prime}\end{bmatrix}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}$. (6)
Translation: Shifts in width and height up to $\pm 10\%$ of image dimensions:
$\begin{bmatrix}x^{\prime}\\ y^{\prime}\end{bmatrix}=\begin{bmatrix}1&0&t_{x}\\ 0&1&t_{y}\end{bmatrix}\begin{bmatrix}x\\ y\\ 1\end{bmatrix}$. (7)
Zoom: Scaling images with factors $s_{x},s_{y}\in[0.8,1.2]$:
$\begin{bmatrix}x^{\prime}\\ y^{\prime}\end{bmatrix}=\begin{bmatrix}s_{x}&0\\ 0&s_{y}\end{bmatrix}\begin{bmatrix}x\\ y\end{bmatrix}$. (8)
Brightness Adjustment: Pixel intensities scaled by $\alpha\in[0.8,1.2]$:
$I^{\prime}(x,y)=\alpha\cdot I(x,y).$ (9)
With these preprocessing and augmentation methods in place, the model learned to handle variations in orientation, scale, position, lighting, and background clutter. This led to reliable recognition in real-world banknote situations.
III-C Model Architecture
@@
III-D Dual-Branch Feature Extraction
III-E Feature Fusion Pipeline
III-F Feature Processing Pipeline
III-F1 L2 Normalization
III-F2 Principal Component Analysis (PCA)
$\mathbf{X}_{\text{PCA}}$ $=\mathbf{X}\cdot\mathbf{V}_{k}$
III-G Compact Multilayer Perceptron Classifier
TABLE III: MLP Classifier Architecture Specifications
Layer Input Output Activation Dropout Parameters
Layer 1 $k$ 512 ReLU 0.5 $(k+1)\times 512$
Layer 2 512 256 ReLU 0.4 131,328
Layer 3 256 128 ReLU 0.3 32,896
Layer 4 128 64 ReLU 0.2 8,256
Layer 5 64 $C$ Softmax $65\times C$
Total MLP Parameters: $\sim$327,665*
*For $k=300$ and $C=9$ currency classes
@@ III-G Compact Multilayer Perceptron Classifier
0: Final Model $\mathcal{M}^{*}$, Performance metrics $\Psi$, Explainability analysis $\{\mathcal{E}_{LIME},\boldsymbol{\Phi}_{SHAP}\}$
@@
19:  Train: $\mathcal{M}_{k}^{*}\leftarrow\arg\min_{\boldsymbol{\theta}}\mathbb{E}[\mathcal{L}_{CCE}(\mathcal{M}_{k}(\mathbf{Z}_{tr}^{(k)}),\mathbf{y}_{tr}^{(k)})]$
@@
40:return $\mathcal{M}^{*}$, $\{\mu_{CV}\pm\sigma_{CV},\Psi,\text{ROC}\}$, $\{\mathcal{E}_{LIME},\boldsymbol{\Phi}_{SHAP}\}$
III-H Evaluation Metrics
@@ III-H1 Accuracy
III-H2 Precision
Precision is the indication of the proportion of true positives among all predicted positives made by the model.
Mathematically, it is expressed by:
@@ III-H4 F1-score
III-H5 Cohen Kappa Coefficient
@@
where $p_{o}$ is the observed agreement among raters and $p_{e}$ is the hypothetical probability of chance agreement, which is calculated using observed data. A value of $\kappa$ > 0.80 is indicated as strong agreement.
III-H6 Matthews Correlation Coefficient
@@ III-H9 Explainable AI methods
where $F$ is the set of all features and $S$ is a subset of features not including $i$
III-I Experimental setting
@@ III-I5 Optimizer
III-I6 Model Complexity and Training Time
The total number of trainable parameters of the proposed model is approximately $5.2$ million, which is significantly lower compared to traditional CNN models such as ResNet50 (23.523.5M parameters). The Tesla T4 GPU had an average training time per epoch of $10$ seconds, and the model was able to converge in 40 or 50 epochs. This shows not only computational efficiency, but also its applicability to real-time applications on resource-constrained and embedded systems.

### v2

@@ I Introduction
  • We have proposed a novel hybrid model that combines MobileNetV3-Large and EfficientNetB0 as complementary feature extractors, followed by a multi-layer perceptron (MLP) pipeline for classification.
@@ III-A1 Data Collection
Data collection is the first and fundamental step of research. To build our Bangladeshi banknote datasets, we have collected images of the currently available and frequently used Bangladeshi currency in 9 denominations (2 Tk, 5 Tk, 10 Tk, 20 Tk, 50 Tk, 100 Tk, 200 Tk, 500 Tk, and 1000 Tk). The images were collected by taking pictures from a mobile camera, a laptop webcam, and from publicly available datasets [29], [30], [40]. Figure 2 shows representative samples from the datasets that illustrate the progressive increase in visual variability and environmental complexity. Formally, the complete dataset $D$ can be expressed as the union of the sources:
@@ III-A2 Dataset Description
(a) Primary Dataset-1
(b) Primary Dataset-2
(c) Custom Dataset-2
(d) Custom Dataset-3
@@
III-B Model Architecture
@@
0: Final Model $\mathcal{M}^{*}$, Performance metrics $\Psi$, Explainability analysis $\{\mathcal{E}_{LIME},\bm{\Phi}_{SHAP}\}$
@@
19:  Train: $\mathcal{M}_{k}^{*}\leftarrow\arg\min_{\bm{\theta}}\mathbb{E}[\mathcal{L}_{CCE}(\mathcal{M}_{k}(\mathbf{Z}_{tr}^{(k)}),\mathbf{y}_{tr}^{(k)})]$
@@
40:return $\mathcal{M}^{*}$, $\{\mu_{CV}\pm\sigma_{CV},\Psi,\text{ROC}\}$, $\{\mathcal{E}_{LIME},\bm{\Phi}_{SHAP}\}$
III-C Dual-Branch Feature Extraction
III-D Feature Fusion Pipeline
III-E Feature Processing Pipeline
III-E1 L2 Normalization
III-E2 Principal Component Analysis (PCA)
$\mathbf{X}_{\text{PCA}}$ $=\mathbf{X}\cdot\mathbf{V}_{k}$
III-F Compact Multilayer Perceptron Classifier
III-G Evaluation Metrics
@@ III-G1 Accuracy
III-G2 Precision
Precision is the indication of the proportion of true positives among all predicted positives made by the model. Mathematically, it is expressed by:
@@ III-G4 F1-score
III-G5 Cohen Kappa Coefficient
@@
where $p_{o}$ is the observed agreement among raters and $p_{e}$ is the hypothetical probability of chance agreement, which is calculated using observed data. A value of $\kappa$ ¿ 0.80 is indicated as strong agreement.
III-G6 Matthews Correlation Coefficient
@@ III-G9 Explainable AI methods
where $F$ is the set of all features and $S$ is a subset of features not including $i$.
TABLE III: MLP Classifier Architecture Specifications
Layer Input Output Activation Dropout Parameters
Layer 1 $k$ 512 ReLU 0.5 $(k+1)\times 512$
Layer 2 512 256 ReLU 0.4 131,328
Layer 3 256 128 ReLU 0.3 32,896
Layer 4 128 64 ReLU 0.2 8,256
Layer 5 64 $C$ Softmax $65\times C$
Total Parameters $\sim$327,665*
III-H Experimental setting
@@ III-H5 Optimizer
III-H6 Model Complexity and Training Time
The total number of trainable parameters of the proposed model is approximately $5.2$ million, which is significantly lower compared to traditional CNN models such as ResNet50 (23.523.5M parameters). The Tesla T4 GPU had an average training time per epoch of $10$ seconds, and the model converged in 40 or 50 epochs. This shows not only computational efficiency, but also its applicability to real-time applications on resource-constrained and embedded systems.